OpenStudy (anonymous):

Which is not a factor of 48? A. 8 B. 14 C. 16 D. 24

OpenStudy (mathlegend):

Which number does not go into 48?

OpenStudy (the_fizicx99):

48/8 = ? 48/14 = ? 48/16 = ? 48/24 = ?

OpenStudy (anonymous):

*facepalm*

OpenStudy (♪chibiterasu):

3 of these numbers share the factor of 8. One does not.

OpenStudy (the_fizicx99):

They all give you a whole number, one gives you a never ending decimal. It's an easier way of seeing which doesn't work.

OpenStudy (♪chibiterasu):

Yeh, process of elimination. ;P Easy way out if you have a calculator on you, or you just know your times tables. :3

OpenStudy (anonymous):

@dreamer02

OpenStudy (anonymous):

@xo_kansasprincess_xo

OpenStudy (anonymous):

Why am I tagged in this..? I only ask for help in math. It's not my best subject. Sorry

OpenStudy (anonymous):

.............

OpenStudy (the_fizicx99):

We basically gave you the answer, if you applied a tiny fraction of effort instead of tagging people to give you the answer you would have already solved this